# From Point Data to Geographic Boundaries: Regionalizing Crowdsourced Latency Measurements

### v3

@@ 1. Introduction
Measuring the performance of Internet access networks is critical for characterizing the quality of service that ISPs deliver to users (Clark and Wedeman, 2021) and for identifying discrepancies in Internet performance in both urban and rural areas (Jiang et al., 2023; Sommers and Barford, 2012). Over the past few decades, there have been significant advancements in measuring access network performance, both in terms of novel measurement infrastructure (Huang et al., 2008; Sundaresan et al., 2014) and analysis techniques (Sundaresan et al., 2011, 2017; Bischof et al., 2017; Saxon and Black, 2022). These advances have primarily been focused on measuring the performance of a single access link, using metrics such as throughput, latency, jitter, and packet loss. The Measurement Lab (M-Lab) (Lab, 2024) and Ookla Speedtest (Ookla, 2024) datasets are currently widely used for understanding Internet performance of an access link (Coalition, 2022). Their increasing prevalence has also enabled researchers to use these datasets to ask a broader set of questions about Internet performance across an ISP or a region (Federal Communications Commission, 2022; New York State Office of the Attorney General, 2020; Battle For the Net, 2022; Pennsylvania State University and Measurement Lab, 2019). In particular, there has been significant recent interest in understanding the distribution of Internet performance across different geographies (Lee et al., 2023; Paul et al., 2023; Sharma et al., 2023a; Clark and Wedeman, 2021), especially for specific social and policy-related inquiries.
@@
Yet, utilizing these crowdsourced measurements to characterize Internet performance across a geographic region introduces new challenges, given the nature of their measurements. Most notably, crowdsourced datasets are self-selected, both in time and in space. Both the Ookla and M-Lab datasets generate crowdsourced “point” measurements from a subset of Internet users across different geographies. These measurements, often irregularly concentrated over space, are performed whenever a user decides to run a speed test. Consequently, these data points reflect only a small, non-uniform fragment of the overall user base and geographic area, posing significant challenges for a comprehensive analysis. Yet, both policy and operational decisions affecting a geographic region rely on applying robust spatial analysis techniques to these small, self-selected samples to make generalizations about Internet performance for the entire resident population across that respective region.
@@
Spatial analysis can potentially transform these scattered data points into cohesive insights, identifying patterns and trends that are not immediately apparent. One significant challenge to this approach is identifying geographic sampling boundaries for Internet performance and determining methods to summarize these point measurements over space. Additionally, individual measurements can be significantly noisy due to various factors such as testing infrastructure, access media, and the client’s hardware or software platform (Paul et al., 2022b; MacMillan et al., 2023). Spatial de-noising and aggregation of these measurements is, therefore, critical for drawing meaningful conclusions about network performance over specific geographies. Analysis based on such de-noised datasets, on the other hand, can ultimately help pinpoint areas needing further infrastructure investments.
@@
Prior work on applying spatial analysis to Internet measurements has taken a different approach, treating spatial boundaries as given and applying aggregation techniques within these pre-defined boundaries. For example, previous work has attempted to characterize the Internet performance over conventional boundaries such as zip codes, census tracts, or neighborhood units (Paul et al., 2021; Saxon and Black, 2022; Sharma et al., 2022; Fiduccia, 2022; Jiang et al., 2023). An important conclusion from previous work is that there is often significant variation in Internet performance over these boundaries (Sharma et al., 2022); previous work has suggested that such regions need additional attention or policy intervention. However, these approaches suffer from a few important limitations. First, the use of aggregate measures such as mean, median and inter-quantile range (IQR) (Paul et al., 2021; Sharma et al., 2022) on point measurements may lead to inaccurate conclusions across a region, particularly when the region exhibits high variability or is poorly sampled. For example, measurements in pre-defined regions are often clustered in some portions of the space and dispersed in others (Jiang et al., 2023). Thus, any aggregate measures calculated over irregularly clustered data may overrepresent densely sampled areas.
@@
Second, to our knowledge, no previous work has assessed the accuracy of previous techniques (Fiduccia, 2022; Jiang et al., 2023) in summarizing Internet performance over a pre-defined geography. This may prevent regulators and ISPs from adopting the most effective aggregation techniques for their analyses. Finally, correlating Internet performance with population measures such as median income and population density (Fiduccia, 2022; Saxon and Black, 2022; Paul et al., 2021) using existing social boundaries may not be inappropriate due to imperfect alignment with infrastructure boundaries. A single social boundary may encompass multiple areas with heterogeneous Internet infrastructure, potentially leading to misleading correlations. Simply put, there is no reason to expect that Internet performance should be spatially aggregated along human-defined boundaries that have nothing to do with the Internet itself. This paper is the first to explore and evaluate alternate spatial clustering approaches that more accurately reflect relationships in the underlying Internet measurement data.
@@
  • We develop an end-to-end analysis pipeline to construct measurement-driven boundaries for sampling Internet performance over a large US city to achieve stable sampling boundaries. Our boundaries show consistency across monthly samples drawn from the same dataset, up to an Adjusted Rand Index of 0.59.
@@
  • We find that boundaries constructed from 17-month-long, ISP-specific data samples do not show significant similarity between ISPs. This shows that a one-size-fits-all measurement data sampling approach may not be appropriate for all ISPs.
  • We release our source code for constructing these boundaries, enabling the research community, policymakers, and ISPs to use it in their analyses (Authors, 2024).
@@ 2.2. Spatial Interpolation and Clustering of Crowdsourced Measurements
In the context of crowdsourced network data, we argue that the noise introduced by factors such as Wi-Fi (Sharma et al., 2023b) and access equipment (Sundaresan et al., 2011) may weaken spatial auto-correlation between neighboring measurements. This can make deterministic techniques more suitable, as they do not rely on the statistical properties of the data as much as stochastic techniques. In this paper, we thus use three deterministic techniques—IDW, LOESS, and STBKR—for comparison and extension. Exploration of stochastic techniques for interpolating Internet measurement data is a ripe avenue for future work.
@@ 3.3. Analysis Approach
We used an 80-20 split of the dataset to evaluate the chosen techniques across the city’s geography. Due to the dataset’s large size, we fit each model on monthly splits, resulting in 17 different fits per interpolation method parameter choice. Our evaluation ensures that all measurements from a single user appear exclusively in either the training set or the test set to prevent overfitting. Including a user’s measurements in both sets could lead models to learn the user’s latency distribution rather than the underlying spatial patterns. By treating each user independently, we assess how well the models generalize to unsampled users at a given location. Upon selecting a suitable interpolation method, we then predict latency at regularly spaced points across the city to construct a uniform surface model of latency. The grid points for this out-of-sample interpolation step are chosen to be spread across the city at a regular spacing of 50 meters. This choice is made to ensure that the interpolated map was evenly spaced, granular and smooth, which allows an unbiased calculation of latency aggregates such as averages and percentiles.
@@ 4. Interpolation
Locally Estimated Scatterplot Smoothing (LOESS) Fits local regression lines to de-noise latency across space. Uses lat-long values directly for regression. $span$ (Proportion of data points for regression)
@@ 6.3. Impact of ISP
When using a collective sample, the aggregation process smoothens out variability in latency measurements across ISPs, leading to more consistent boundaries over time. The differences between ISPs may arise from several factors such as the geographic distribution of servers, the underlying access technology, the network design for the ISPs, or the per-ISP sample size. When measurements from all ISPs are jointly interpolated, individual ISP characteristics are more likely to be smoothened out, and the resulting boundaries are more likely to reflect the geographic distribution of latency in the region. Therefore, a one-size-fits-all approach to sampling boundaries may not be suitable for all ISPs with the current method.
@@ 9. Conclusion & Future Work
Our work provides a method for identifying Internet latency sampling boundaries, assuming minimal information about the underlying infrastructure. By delineating clear latency boundaries, network operators can target specific areas for infrastructure improvements, optimizing resource allocation and enhancing overall network performance. Policymakers can apply these insights to develop regulations and incentives that promote equitable access to low latency internet across different regions, and addressing the digital divide. Although our approach helps reveal significant spatial structure from latency measurements, our sampling boundaries are still based on crowdsourced data. Despite of its wide adoption, Ookla data may not be representative of all users, especially those with limited or no Internet access. Future work can thus explore the possibility of exploring an augmentation of multiple data sources using novel metrics to improve the representativeness of the data. Our findings from Section 6.1 provide a starting point for such an exploration. The design of spatial clustering methods that alleviate the ARI scores for metrics such as the inequality ratio, or latency reduction, can be a promising direction. Validating these boundaries using ground truth data from network operators and regulators to ensure their accuracy and reliability is another important future direction.
@@
A key component of our method relies on prior interpolation methods to build a surface model of latency across the city. Although we primarily use IDW interpolation in this work, it remains to be seen how much boundaries produced by different interpolation methods agree. Additionally, interpolation methods that incorporate local context such as network topology, routing information, and urban infrastructure can be explored to improve the accuracy of the surface model. This information can enhance the stability of the clustering boundaries and the accuracy of the regionalization process.
@@
The primary focus of our analysis is on the city of Chicago. While Chicago’s diverse urban environment offers an intriguing testbed, we expect our methodology to be generalizable to other cities with similar data availability and population density. A key challenge to extending our work to rural and remote areas with low connectivity is the availability of crowdsourced data. Operators and regulators can play an important role in collecting and sharing data from these regions to explore the potential of our methodology in these areas.

### v4

# Beyond Data Points: Regionalizing Crowdsourced Latency Measurements

@@ Abstract.
Access networks, Latency, Crowdsourcing, Interpolation, Spatial analysis
ccs: Networks Network performance analysisccs: Networks Network measurementcopyright: rightsretainedjournal: POMACSjournalyear: 2024journalvolume: 8journalnumber: 3article: 34publicationmonth: 12doi: 10.1145/3700416
@@ 1. Introduction
Measuring the performance of Internet access networks is critical for characterizing the quality of service that ISPs deliver to users (Clark and Wedeman, 2021) and for identifying discrepancies in Internet performance in both urban and rural areas (Jiang et al., 2023; Sommers and Barford, 2012). Over the past few decades, there have been significant advancements in measuring access network performance, both in terms of novel measurement infrastructure (Huang et al., 2008; Sundaresan et al., 2014) and analysis techniques (Sundaresan et al., 2011, 2017; Bischof et al., 2017; Saxon and Black, 2022). These advances have primarily   focused on measuring the performance of a single access link, using metrics such as throughput, latency, jitter, and packet loss. The Measurement Lab (M-Lab) (Lab, 2024) and Ookla Speedtest (Ookla, 2024) datasets are currently widely used for understanding Internet performance of an access link (Coalition, 2022). Their increasing prevalence has also enabled researchers to use these datasets to ask a broader set of questions about Internet performance across an ISP or a region (Federal Communications Commission, 2022; New York State Office of the Attorney General, 2020; Battle For the Net, 2022; Pennsylvania State University and Measurement Lab, 2019). In particular, there has been significant recent interest in understanding the distribution of Internet performance across different geographies (Lee et al., 2023; Paul et al., 2023; Sharma et al., 2023a; Clark and Wedeman, 2021), especially for specific social and policy-related inquiries.
@@
Yet, utilizing these crowdsourced measurements to characterize Internet performance across a geographic region introduces new challenges, given the nature of their data. Most notably, crowdsourced datasets are self-selected, both in time and in space. Both the Ookla and M-Lab datasets generate crowdsourced “point” measurements from a subset of Internet users across different geographies. These measurements, often irregularly concentrated over space, are performed whenever a user decides to run a speed test. Consequently, these data points reflect only a small, non-uniform fragment of the overall user base and geographic area, posing significant challenges for a comprehensive analysis. Yet, both policy and operational decisions affecting a geographic region rely on applying robust spatial analysis techniques to these small, self-selected samples to make generalizations about Internet performance for the entire resident population across that respective region.
@@
Spatial analysis can potentially transform these scattered data points into cohesive insights, identifying patterns and trends that are not immediately apparent. One significant challenge to this approach is identifying geographic sampling boundaries for Internet performance and determining methods to summarize these point measurements over space. Additionally, individual measurements can be significantly noisy due to various factors such as testing infrastructure, access media, and the client’s hardware or software platform (Paul et al., 2022b; MacMillan et al., 2023). Spatial de-noising and aggregation of these measurements is, therefore, critical for drawing meaningful conclusions about network performance over specific geographies. Analysis based on such de-noised datasets, on the other hand, can ultimately help pinpoint areas truly needing further infrastructure investments.
@@
Prior work on applying spatial analysis to Internet measurements has taken a different approach, treating spatial boundaries as given and applying aggregation techniques within these pre-defined boundaries. For example, previous work has attempted to characterize Internet performance over conventional boundaries such as zip codes, census tracts, or neighborhood units (Paul et al., 2021; Saxon and Black, 2022; Sharma et al., 2022; Fiduccia, 2022; Jiang et al., 2023). An important conclusion from previous work is that there is often significant variation in Internet performance over these boundaries (Sharma et al., 2022); previous work has suggested that such regions need additional attention or policy intervention. However, these approaches suffer from a few important limitations. First, the use of aggregate measures such as mean, median and inter-quantile range (IQR) (Paul et al., 2021; Sharma et al., 2022) on point measurements may lead to inaccurate conclusions across a region, particularly when the region exhibits high variability or is poorly sampled. For example, measurements in pre-defined regions are often clustered in some portions of the space and dispersed in others (Jiang et al., 2023). Thus, any aggregate measures calculated over irregularly clustered data may overrepresent densely sampled areas.
@@
Second, to our knowledge, no previous work has assessed the accuracy of previous techniques (Fiduccia, 2022; Jiang et al., 2023) in summarizing Internet performance over a pre-defined geography. This may prevent regulators and ISPs from adopting the most effective aggregation techniques for their analyses. Finally, correlating Internet performance with population measures such as median income and population density (Fiduccia, 2022; Saxon and Black, 2022; Paul et al., 2021) using existing social boundaries may be inappropriate due to imperfect alignment with infrastructure boundaries. A single social boundary may encompass multiple areas with heterogeneous Internet infrastructure, potentially leading to misleading correlations. Simply put, there is no reason to expect that Internet performance should be spatially aggregated along human-defined boundaries that have nothing to do with the Internet itself. This paper is the first to explore and evaluate alternate spatial clustering approaches that more accurately reflect relationships in the underlying Internet measurement data.
@@
  • We develop an end-to-end analysis pipeline to construct stable measurement-driven boundaries for sampling Internet performance over a large US city . Our boundaries show consistency across monthly samples drawn from the same dataset, up to an Adjusted Rand Index of 0.59.
@@
  • We find that boundaries constructed from 17-month-long, ISP-specific data samples do not show significant similarity between ISPs. This suggests that the FCC should consider releasing ISP-specific representations of Internet performance for greater transparency.
  • While network operators may deploy their own measurement infrastructures, our approach offers significant advantages by utilizing crowdsourced data, allowing coverage from multiple real-world vantage points in a cost-effective manner. We release our source code for constructing these boundaries, enabling the research community, policymakers, and ISPs to use it in their analyses (Sharma et al., 2024).
@@ 2.2. Spatial Interpolation and Clustering of Crowdsourced Measurements
In the context of crowdsourced network data, we argue that the noise introduced by factors such as Wi-Fi (Sharma et al., 2023b) and access equipment (Sundaresan et al., 2011) may weaken spatial auto-correlation between neighboring measurements. This can make deterministic techniques more suitable, as they do not rely on the statistical properties of the data as much as stochastic techniques. In this paper, we thus use three deterministic techniques—IDW, LOESS, and STBKR—for comparison and design an analysis pipeline that can be integrated with any spatial interpolation method. Exploration of stochastic techniques for interpolating Internet measurement data is a ripe avenue for future work.
@@ 3.1. Scope of Analysis
 [^1] Access to the Ookla dataset used in this paper can be obtained from https://www.ookla.com/speedtest-intelligence
@@ 3.2. Data Preprocessing
# hexagonal cells 899
# cells with lower than average sample size (sparse cells) 643
Average # measurements per cell per month 66.15
@@ 3.3. Analysis Approach
We used an 80-20 split of the dataset to evaluate the chosen techniques across the city’s geography. Due to the dataset’s large size, we fit each model on monthly splits, resulting in 17 different fits per interpolation method parameter choice. Our evaluation ensures that all measurements from a single user appear exclusively in either the training set or the test set to prevent overfitting. Including a user’s measurements in both sets could lead models to learn the user’s latency distribution rather than the underlying spatial patterns. By treating each user independently, we assess how well the models generalize to unsampled users at a given location. Upon selecting a suitable interpolation method, we then predict latency at regularly spaced points across the city to construct a uniform surface model of latency. The grid points for this out-of-sample interpolation step are chosen to be spread across the city at a regular spacing of 50 meters. This choice is made to ensure that the interpolated map was evenly spaced, granular and smooth, which allows an unbiased calculation of latency aggregates such as averages and percentiles. We implement our interpolation workflow using standard Python libraries such as Scikit-learn (Pedregosa et al., 2011) and Geopandas (Jordahl et al., 2020), but tools such as ArcGIS (Redlands, 2011) or QGIS (QGIS Development Team, 2009) can also be adopted for this step. The choice of Python libraries is motivated by their open-source nature, flexibility, and ease of integration with typical data science workflows.
@@ 4. Interpolation
Locally Estimated Scatterplot Smoothing (LOESS) Fits local regression lines to de-noise latency across space. Uses lat-long values directly for regression. $span$ (Proportion of data points used for regression)
@@ 6.3. Impact of ISP
When using a collective sample, the aggregation process smoothens out variability in latency measurements across ISPs, leading to more consistent boundaries over time. The differences between ISPs may arise from several factors such as the geographic distribution of servers, the underlying access technology, the network design for the ISPs, or the per-ISP sample size. When measurements from all ISPs are jointly interpolated, individual ISP characteristics are more likely to be smoothened out, and the resulting boundaries are more likely to reflect the geographic distribution of latency in the region. Looking at each ISP individually on a regionalized map is likely to provide a more consumer-transparent view of network performance in a region. It would not only help new subscribers make informed decisions about their Internet connections, but also assist in maintaining greater scrutiny over ISPs regarding their service level agreements (SLAs). The FCC should thus consider releasing ISP and region specific maps to ensure that the representations of network performance are accurate and reliable.
@@ 9. Conclusion & Future Work
Our work provides a method for identifying Internet latency sampling boundaries, assuming minimal information about the underlying infrastructure. By delineating clear latency boundaries, network operators can target specific areas for infrastructure improvements, optimizing resource allocation and enhancing overall network performance. For instance, our illustration from Figure 3(b) provides a starting point for identifying areas with higher latency. Given a sample of diagnostic measurement data and a specific timescale, contiguous regions with poor latency can be identified over a geography in a similar manner. Even though network operators may have their own measurement infrastructure in place, our method is particularly valuable for leveraging crowdsourced data, which can provide coverage from a large number of real-world vantage points in a cost-effective manner. It enables operators to work with scattered samples, where continuous measurements are unavailable. Integrating our clustering approach with further diagnostic information can give rise to suitable infrastructure improvements such as cable upgrades, deployment of additional hardware, or routing optimizations in under-provisioned areas. Our finding from Section 6.2 suggests that the use of regular spatial units in place of administrative boundaries is expected to generate Internet performance representations that are more appropriate to inform long-term policy interventions. Finally, our finding from Section 6.3 provides an important insight into using crowdsourced data for the FCC. Instead of relying on a large heterogeneous dataset, regulators can use our method to identify boundaries on a per-ISP basis, which can help in understanding the impact of different ISPs on overall network performance in a region.
Reliance on crowdsourced measurement data.
Although our approach helps reveal significant spatial structure from latency measurements, our sampling boundaries are still based on crowdsourced data. Despite its wide adoption, Ookla data may not be representative of all users, especially those with limited or no Internet access. Future work can thus explore the possibility of exploring an augmentation of multiple data sources using novel metrics to improve the representativeness of the data. Our findings from Section 6.1 provide a starting point for such an exploration. The design of spatial clustering methods that improve the ARI scores for metrics such as the inequality ratio, or latency reduction, can be a promising direction, as these metrics can potentially remain consistent across datasets and device types.
@@
A key component of our method relies on prior interpolation methods to build a surface model of latency across the city. Although we primarily use IDW interpolation in this work, the mutual agreement between boundaries drawn using different interpolation algorithms is remaining to be evaluated. Additionally, interpolation methods that incorporate local context such as network topology, routing information, and urban infrastructure can be explored to improve the accuracy of the surface model. This information can enhance the stability of the clustering boundaries and the accuracy of the regionalization process.
@@
The primary focus of our analysis is on the city of Chicago. Our approach identifies boundaries for sampling Internet latency without relying on Chicago-specific information such as census or demographic data. While Chicago’s diverse urban environment offers an intriguing testbed, we expect our methodology to be generalizable to other cities with similar data availability and population density. A key challenge to extending our work to rural and remote areas with low connectivity is the availability of crowdsourced data. Operators and regulators can play an important role in collecting and sharing data from these regions to explore the potential of our methodology in these areas.
@@
Acknowledgments
We thank the anonymous reviewers and our shepherd Paul Barford for their constructive feedback and guidance. We are grateful to Jared Schachner, Jonatas Marques, and other colleagues at the Internet Equity Initiative (IEI) for their feedback on this research. This work was supported by National Science Foundation (NSF) grants IMR-2319603, CCRI-2213821, and SAI-2324515 at the University of Chicago, and by the ANR Project No ANR-21-CE94-0001-01 (MINT) at ENS de Lyon.